OpenStudy (avanti):

PRECALC QUESTIONS: 1. Find 2 common angles that sum to (17pi/12) 2. Evaluate tan(17pi/12) using the sum identity for tangent. Please help, im very confused!!

OpenStudy (anonymous):

We want to express `(17pi)/12` as a sum/difference of common angles. The angles usually memorized are the quadrantal angles (`0,pi/2,pi,(3pi)/2,2pi,` etc...) and the multiples of `pi/6,pi/4,pi/3` . There are an infinite number of possible answers. Since the denominator is 12 we might select a sum/difference of angles with denominators 3 and 4. Then `(17pi)/12=(Api)/3+(Bpi)/4` . Adding the fractions on the right hand side together we get `(4Api+3Bpi)/12` . Then setting the numerators equal we get `4A+3B=17` which again has an infinite number of solutions. One solution is to let A=2 and B=3. Thus `(17pi)/12=(2pi)/3+(3pi)/4`

OpenStudy (anonymous):

I realized lol...

OpenStudy (dumbcow):

you could also use \[\frac{17\pi}{12} = \frac{5\pi}{4}+\frac{\pi}{6}\]

OpenStudy (avanti):

okay thanks! how do i do the second part?

OpenStudy (dumbcow):

\[\tan(a+b) = \frac{\tan a + \tan b}{1-\tan a \tan b}\]

OpenStudy (avanti):

how do i know what values a and b are?

OpenStudy (anonymous):

17pi is a and 12 is b

OpenStudy (anonymous):

So exactly how the original problem reads

OpenStudy (avanti):

so its tan(17pi + 12) = tan(17pi) + tan(12) / 1-tan(17pi)tan(12) ?

OpenStudy (dumbcow):

??

OpenStudy (dumbcow):

\[\tan(\frac{17\pi}{12}) = \tan(\frac{2\pi}{3}+\frac{3\pi}{4}) = \tan(\frac{\pi}{6}+\frac{5\pi}{4})\]

OpenStudy (dumbcow):

?? looks like you need to review your basic trig you are asked to evaluate tan(17pi/12) without calculator the only way to do this is use common angles for which tangent ratios are known use the formula above for angle addition to obtain exact answer for tan(17pi/12) notice you have 2 possible combinations of common angles to use...either way you will get same answer

OpenStudy (avanti):

i know that you need to add the 2 angles which i did to get 255 degrees (and then of course take the tangent of that) but what i don't understand is how the sum identity has anything to do with solving it.

OpenStudy (dumbcow):

well 17pi/12 = 255 ....so adding the angles to get 255 is unnecessary the question is how do you intend to find tan(255) without calculator? ....hence the sum identity is needed:)

OpenStudy (avanti):

okay thank you for all your help :)

OpenStudy (dumbcow):

yw

OpenStudy (dumbcow):

here is solution for you to check your work http://www.wolframalpha.com/input/?i=tan%2817pi%2F12%29

OpenStudy (avanti):

cool thanks again!